# Feasibility Analysis of In-Band Coexistence in Dense LEO Satellite Communication Systems

### 01 Dec 2023 v2

@@ VI System Performance under Uncertainty
In Figure 14, we highlight how the level of uncertainty $\gamma$ and the interference threshold $\mathsf{INR}_{\mathrm{th}}$ impact the number of feasible secondary satellites $N_{{\mathbf{s}}}^{\prime}$. First, in Figure 14a, we depict the empirical CDF of $N_{{\mathbf{s}}}^{\prime}$ over time for various $\gamma$ when the threshold is $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB. Notice that the shape of the distribution remains fairly unchanged but undergoes a shift leftward as uncertainty $\gamma$ increases. In median, there are typically more than $4$ fewer feasible satellites when $\gamma=40^{\circ}$, compared to when ${{\mathbf{p}}}^{\star}$ is known exactly. In Figure 14b, we depict a heatmap of the time-averaged $N_{{\mathbf{s}}}^{\prime}$ for a wide range of $\gamma$ and $\mathsf{INR}_{\mathrm{th}}$. Even with extreme uncertainty and a very strict threshold, on average there are $3$ satellites that can guarantee the protection constraint is met across the entire set of possible primary satellites $\mathcal{P}^{\prime}$. At $\mathsf{INR}_{\mathrm{th}}=-12$ dB, there are more than 10 feasible satellites on average when uncertainty is $40^{\circ}$ or less. With at most $25^{\circ}$ of uncertainty, the average number of feasible satellites decreases from 17 to 12, compared to a virtually unconstrained and perfectly known scenario (i.e., when $\mathsf{INR}_{\mathrm{th}}=0$ dB and $\gamma=0^{\circ}$).
@@
In Figure 15a, we show the distribution of guaranteed SINR achieved by the secondary system over time, normalized to its upper bound, the maximum SNR (i.e., $\mathsf{SNR}({\mathtt{v}},{{\mathbf{s}}}_{\infty}^{\dagger})$ from (17)), where $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB. Without any uncertainty $\gamma=0^{\circ}$, there is at worst about a $1$ dB gap from its upper bound, which is less than $0.25$ dB the overwhelming majority of the time. However, when ${{\mathbf{p}}}^{\star}$ is not known precisely, uncertainty leads to more severe losses. For $\gamma=20^{\circ}$, the guaranteed SINR has a more substantial lower tail, with a gap of more than a $2$ dB about 10% of the time; this grows to over 80% when $\gamma=40^{\circ}$. Nonetheless, it is a welcome sight to observe that at most around $4$ dB is sacrificed by the secondary system in protecting (and being interfered by) the primary system. Figure 15b extends this analysis by depicting the median guaranteed secondary SINR for various interference thresholds and levels of uncertainty. Immediately, it can be observed in the most extreme case that the minimum guaranteed SINR is above $3.5$ dB in median, about $2$ dB short of the upper bound. In terms of median guaranteed SINR, the secondary system is not significantly impacted by making the protection threshold more strict under modest uncertainty. When uncertainty exceeds $40^{\circ}$, however, we begin to see that more stringent $\mathsf{INR}_{\mathrm{th}}$ leads to more dramatic losses in median guaranteed SINR.
@@
From Figure 15a, it is certainly remarkable at first glance that the secondary system sacrifices at most $4$ dB in guaranteed SINR under an uncertainty of $\gamma=40^{\circ}$ and a strict interference constraint of $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB. Given the system parameters and employed methodology outlined in Section III, the SNR delivered by a secondary satellite ${{\mathbf{s}}}\in\mathcal{S}$ is dictated mostly by the path loss from the satellite ${{\mathbf{s}}}$ and ground user ${\mathtt{v}}$. This path loss depends on the altitude and elevation angle of the satellite [39], and we typically observe about a $4$ dB difference in delivered SNR across overhead satellites, with their altitudes as in Table II and $\epsilon_{\mathrm{min}}=35^{\circ}$. Combining this with a strict interference constraint results in a roughly $4$ dB difference in SINR. This, along with the fact that we see multiple feasible secondary satellites under uncertainty, leads to the selection of one which sacrifices at most about $4$ dB in SINR.
@@ Acknowledgments
We thank Arun Ghosh, Anil Rao, and Vikram Chandrasekhar from Amazon’s Project Kuiper for their valuable discussions and feedback on our methodology and results.

### 10 Dec 2023 v3

@@ VI System Performance under Uncertainty
In Figure 14, we highlight how the level of uncertainty $\gamma$ and the interference threshold $\mathsf{INR}_{\mathrm{th}}$ impact the number of feasible secondary satellites $N_{{\mathbf{s}}}^{\prime}$. First, in Figure 14a, we depict the empirical CDF of $N_{{\mathbf{s}}}^{\prime}$ over time for various $\gamma$ when the threshold is $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB. Notice that the shape of the distribution remains fairly unchanged but undergoes a shift leftward as uncertainty $\gamma$ increases. In median, there are typically more than $4$ fewer feasible satellites when $\gamma=40^{\circ}$, compared to when ${{\mathbf{p}}}^{\star}$ is known exactly, and 8 fewer when $\gamma=50^{\circ}$. In Figure 14b, we depict a heatmap of the time-averaged $N_{{\mathbf{s}}}^{\prime}$ for a wide range of $\gamma$ and $\mathsf{INR}_{\mathrm{th}}$. Even with extreme uncertainty and a very strict threshold, on average there are $3$ satellites that can guarantee the protection constraint is met across the entire set of possible primary satellites $\mathcal{P}^{\prime}$. At $\mathsf{INR}_{\mathrm{th}}=-12$ dB, there are more than 10 feasible satellites on average when uncertainty is $40^{\circ}$ or less. With at most $25^{\circ}$ of uncertainty, the average number of feasible satellites decreases from 17 to 12, compared to a virtually unconstrained and perfectly known scenario (i.e., when $\mathsf{INR}_{\mathrm{th}}=0$ dB and $\gamma=0^{\circ}$).
@@
In Figure 15a, we show the distribution of guaranteed SINR achieved by the secondary system over time, normalized to its upper bound, the maximum SNR (i.e., $\mathsf{SNR}({\mathtt{v}},{{\mathbf{s}}}_{\infty}^{\dagger})$ from (17)), where $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB. By normalizing the guaranteed SINR in this way, we can measure the degradation in secondary system signal quality due to the combined effects of (i) abiding by the protection constraint under uncertainty and (ii) interference from the primary system. Without any uncertainty $\gamma=0^{\circ}$, there is at worst about a $2.5$ dB gap in guaranteed SINR from its upper bound and less than a $1$ dB gap the majority of the time. When ${{\mathbf{p}}}^{\star}$ is not known precisely, this uncertainty leads to losses in guaranteed SINR, especially at the lower tails, but this is most apparent only under high uncertainty. For $\gamma\leq 20^{\circ}$, the guaranteed secondary SINR falls short by only a fraction of a dB from that with perfect knowledge of ${{\mathbf{p}}}^{\star}$. When uncertainty increases to $\gamma=50^{\circ}$, there is only about a $1$ dB sacrifice in median, but a much more substantial lower tail is present. Nonetheless, it is a welcome sight that, the overwhelming majority of the time, less than $4$ dB is sacrificed by the secondary system in protecting (and being interfered by) the primary system, even with very limited knowledge on the primary serving satellite.
Figure 15b extends this analysis by depicting the median guaranteed secondary SINR (unnormalized) for various interference thresholds and levels of uncertainty. It can be observed in the most extreme case that the minimum guaranteed SINR is above $3.5$ dB in median, about $2$ dB short of the upper bound. In terms of median guaranteed SINR, the secondary system is not significantly impacted by making the protection threshold more strict under modest uncertainty. When uncertainty exceeds $40^{\circ}$, however, we begin to see that more stringent $\mathsf{INR}_{\mathrm{th}}$ leads to more dramatic losses in median guaranteed SINR.
@@
From Figure 15a, it is certainly remarkable at first glance that the secondary system sacrifices at most only $4$ dB in guaranteed SINR under an uncertainty of $\gamma=50^{\circ}$ and a strict interference constraint of $\mathsf{INR}_{\mathrm{th}}=-12.2$ dB, compared to exact knowledge of ${{\mathbf{p}}}^{\star}$. The SNR delivered by a secondary satellite ${{\mathbf{s}}}\in\mathcal{S}$ is dictated mostly by its path loss to the ground user, which itself depends on the satellite altitude and elevation angle [39]. Across all overhead satellites, there is typically 44445555 dB of variability in SNR delivered to a given user [36]. Combining this with the fact that satisfying the interference protection constraint often inherently reduces interference onto the secondary ground user, a feasible secondary satellite typically only sacrifices at most roughly 44445555 dB in SINR when $\mathsf{INR}_{\mathrm{th}}$ is strict. Selecting the feasible satellite which maximizes SINR therefore often results in less than 44445555 dB of loss.
@@ Acknowledgments
We thank Arun Ghosh, Michael Hicks, Anil Rao, and Vikram Chandrasekhar from Amazon’s Project Kuiper for their valuable discussions and feedback on our methodology and results.